A wife and husband, Linda and Mark, have ages that total 80 years. Two years ago, 2 times Marks age was 2 years more than, 3 times Lindas age. How old is each now? (Write and solve the equation)
Answer by flame8855(424)

w is the age of the wife
h is the age of the husband
w + h = 80
2(h-2) = 2 + 3 ( w-2)
2h -4 = 2 +3w -6
2h -3w = 0
2h = 3w
2h= 3(80-h)
2h = 240 -3h
5h = 240
h = 48 ( Mark)
w = 32 (Linda)